please help with part D-H!Problem 2 Mr. Lee runs an orange grove, SweetOrange. Its harvest season and each week using labor services (L) and equipment
Answer #1

ƏMP q = 250443 KY3 Page 1 9) MP = 2q = 250 ko * ¢3! = 83.33 ila kis SOT = 83.33 (-3) 1*-1kco So, there are diminishing ret

Page 2 C = WLtok = 100L+100k dil) A = 250 3 k5 + 2 (0-100L-100k) d.2) FOC 24 0,450, R = 0 d.3) 40 = 83.33 (213K!3 - 1001 =

e)

Lowest TC = w*L + r*K

TC = 100 * 8 + 100 * 8

TC = $1600

f)

share represented by labor cost = w*L

= 800

share represented by equipment = r*K

= 800

We observe that both shares are equal as the parameters given in the production function are also equal (1/3)

g)

Average Cost = TC / q

= 1600 / 1000

= 1.6

h)

Since, AC = TC / q

Increase in the denominator from 1000 to 4000 will reduce the average cost

new AC = 1600 / 4000

= 0.4

**if you liked the answer, then please upvote. Would be motivating for me. Thanks